OpenStudy (haleyelizabeth2017):

Determine the equation of the horizontal asymptotes, if any, of the function. f(x)=(2x+1)/(x+1)

OpenStudy (haleyelizabeth2017):

\[f(x)=\frac{ 2x+1 }{ x+1 }\]

OpenStudy (rizags):

are you allowed to use limits?

OpenStudy (haleyelizabeth2017):

I'm not sure.

OpenStudy (rizags):

what class is this for?

OpenStudy (haleyelizabeth2017):

PreCalculus

OpenStudy (rizags):

ok then maybe some limits are allowed

OpenStudy (rizags):

are you familiar with limits?

OpenStudy (haleyelizabeth2017):

Nope. But...the book gives us two methods we can use. Solving for x in terms of y is one method, and divide the numerator and denominator by the highest power of x is the second.

OpenStudy (rizags):

yes use the second

OpenStudy (rizags):

multiply the whole thing by \[\large \frac{\frac{1}{x}}{\frac{1}{x}}\]

OpenStudy (anonymous):

the numerator and denominator are both polynomials the degrees are the same (they are both of degree 1) the horizontal asymptotes is the ratio of the leading coefficients

OpenStudy (anonymous):

in your example it is \[y=\frac{2}{1}=2\]

OpenStudy (anonymous):

it always works this way if the degrees are the same if they are not, then it is different

OpenStudy (haleyelizabeth2017):

Okay, sorry...was afk for a moment there

OpenStudy (anonymous):

if the degree of the denominator is larger, then it is \(y=0\) if the degree of the numerator is larger, then there is no horizontal asymptote

OpenStudy (haleyelizabeth2017):

Okay, good to know. Thank you